OpenStudy (anonymous):

Need Help finding the function value of h(x)=5x-1

OpenStudy (anonymous):

For what value of x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am asking what is the value of x for which you need the function value? That information is missing from the question.

OpenStudy (anonymous):

well it was a question with multiple choice answer like a. h(4) b.h(8) c. h(-3) d.h(-4) e. h(a-1) f. h(a)+3

OpenStudy (anonymous):

It is not multiple choice. It is saying you should find the function value for all those values of x. 4, 8, -3, -4, a-1 etc.... To find the value of h(x) for specific value of x, you simply have to substitute the value of x in the function. h(4) = 5*4 - 1= 20 - 1 = 19 h(8) = 5*8 - 1 = 40 - 1 = 39 And...so on....

OpenStudy (anonymous):

Ah I see that makes it simpler now thank you so much!